Multiply every term in the equation 2x^2 + 10x + 12 = 0 by a number other than 0 that has not been used in this discussion thread already, and solve for x by factoring.
These numbers can not be used: 0, 1, 4, 6,

Answer by stanbon(75887)

Multiply every term in the equation 2x^2 + 10x + 12 = 0 by a number other than 0 that has not been used in this discussion thread already, and solve for x by factoring.
These numbers can not be used: 0, 1, 4, 6,
------
Multiply through by 1/2 to get:
x^2 + 5x + 6 = 0
Factor:
(x+3)(x+2) = 0
x = -3 or x = -2
====================
Cheers,
Stan H.